OpenStudy (anonymous):

Three consecutive terms in the expansion of (1+x)^n have coefficients 120,210, and 252. Find n. BINOMIAL THEOREM QUESTION @ganeshie8

OpenStudy (redheadangel):

First line means (a + b)^0 = 1, then (a + b)^1 = 1·a + 1·b; ...; (a + b)³ = 1·a³ + 3·a²·b + 3·a·b² + 1·b³ etc. Find your coefficients in pascal's triangle.

OpenStudy (redheadangel):

what you have to do is apply the binomial expansion theorem, and figure out which n gives you those three coefficients in that order.